A wire of length 0.3 m, carrying a current of 2 A, is just floating in a magnetic field of flux density

5 × 10^-2 T which is at right angles to the wire. Calculate the mass of the wire.




1
Expert's answer
2021-03-14T19:21:18-0400

In order to wire be floating in a magnetic field, the magnetic force must balance force of gravity:


"F_B=F_g,""BILsin\theta=mg,""m=\dfrac{BILsin\theta}{g},""m=\dfrac{5\cdot10^{-2}\ T\cdot2\ A\cdot0.3\ m\cdot sin90^{\circ}}{9.8\ \dfrac{m}{s^2}}=3.1\cdot10^{-3}\ kg=3.1\ g."